The formula N=2x^2+4x+1 represents the number of households. N, in thousands, in a certain city that have a computer x years after 1990. According to the formula, in what year were there 97 thousand households with computers in this city?
Answer by ewatrrr(24785)

 
Hi,
N=2x^2+4x+1 , where N, in thousands, have a computer x years after 1990.
Solve in what year were there 97 thousand households with computers
97 = 2x^2 + 4x 1
2x^2 + 4x - 96 = 0
x^2 + 2x - 48 = 0
factoring
(x+8)(x-6)= 0
x = 6 |tossing out negative solution for number of years.
1990 + 6 = 1996, the year there were 97 thousand households with computers
2*6^2 + 4*6 + 1 = 72 + 24 + 1 = 97 |checking our answer